Tavon has a 50$ gift card that looses 2$ for each 30 day period it isn’t used.He has a 40$gift card that looses 1.50 for each 30 day period it isn’t used. Write and solve an equation for the number of 30 day periods until the value of the gift cards will be equal.What will the value be of each card be when they have equal value?
1
Expert's answer
2019-08-16T09:48:58-0400

50 – 2x = 40 – 1.5x

10 = 0.5x

x = 20

so after 20 30-day-periods the value of gift cards will be equal

the value on each card will be 10

50 - 2*20 = 10 = 40 - 1.5*20